OpenStudy (anonymous):

. Show all work. A disc jockey has 11 songs to play. Eight are slow songs, and three are fast songs. Each song is to be played only once. In how many ways can the disc jockey play the 11 songs if • The songs can be played in any order. • The first song must be a slow song and the last song must be a slow song. • The first two songs must be fast songs.

OpenStudy (anonymous):

hi polpak how r u today?

OpenStudy (anonymous):

The number of ways you can permute k elements from a set of n elements is \[P(n,k) = \frac{n\!!}{(n-k)\!!}\] \[\implies P(11,11) = \frac{11!}{(11-11)!} = \frac{11!}{0!} = 11!\]

OpenStudy (anonymous):

I'm doing well!

OpenStudy (anonymous):

there is only 11 songs played in order?

OpenStudy (anonymous):

No. There's 11 factorial different ways the jocky can play 11 songs with no restriction on the order.

OpenStudy (anonymous):

= 39916800

OpenStudy (anonymous):

how did u get that number? 11*11=121

OpenStudy (anonymous):

11*10*9*8*7*6*5*4*3*2*1 = 11! = 39916800

OpenStudy (anonymous):

oh ok.. so u take each number 1-11 and multiply them..

OpenStudy (anonymous):

Right. When he is choosing the first song there are 11 different ways he can pick it. After that there are 10 songs he can choose from for the second song. etc.

OpenStudy (anonymous):

ok.. now what about the 2nd question..

OpenStudy (anonymous):

Ok, well how many different songs can he pick from for the first song?

OpenStudy (anonymous):

39916800

OpenStudy (anonymous):

No, there aren't that many songs to choose from. Also he has to pick a slow song first.

OpenStudy (anonymous):

so there is 11 songs total. and 8 r slow so it would be 8 songs to choose from that are slow

OpenStudy (anonymous):

Ok, so there are 8 different choices for the first song. How many options are there for the last song (after the first song gets picked)?

OpenStudy (anonymous):

7 options

OpenStudy (anonymous):

Ok, so he can pick the first and last songs 8*7 = 56 different ways. Once he has done that he just has to figure out the order for the rest of the songs.. How many are there?

OpenStudy (anonymous):

How many songs are left to pick from that is?

OpenStudy (anonymous):

9 songs left.. after the 2 are taken out.. for being slow

OpenStudy (anonymous):

or is it from the 56 ways?

OpenStudy (anonymous):

No that's right. There are 9 songs left you have to pick an order for. So there is 9*8*...*2*1= 9! different ways to order those. Therefore there is in total: 56*9! = 20,321,280 different ways you can play the songs with a slow at the beginning and a slow one at the end.

OpenStudy (anonymous):

ok thats makes sense.

OpenStudy (anonymous):

Now try the last one

OpenStudy (anonymous):

there are 3 songs that are fast

OpenStudy (anonymous):

and the first 2 have to be fast

OpenStudy (anonymous):

2*1=3

OpenStudy (anonymous):

2*1 = 2 actually.

OpenStudy (anonymous):

oh yea.. duh lol

OpenStudy (anonymous):

But the question is... When he goes to pick the first song, how many different ways can he do it?

OpenStudy (anonymous):

2*11 ?

OpenStudy (anonymous):

No. How many songs can he choose from for the first song?

OpenStudy (anonymous):

3 songs to chose from

OpenStudy (anonymous):

cuz they have to be fast songs

OpenStudy (anonymous):

Right. And how many does he have to choose from after he picks the first one?

OpenStudy (anonymous):

For the second song

OpenStudy (anonymous):

2 songs

OpenStudy (anonymous):

cuz the second one has to be fast also

OpenStudy (anonymous):

And how many can he pick from for the 3rd song?

OpenStudy (anonymous):

9 songs for the 3rd

OpenStudy (anonymous):

And for the fourth?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

etc

OpenStudy (anonymous):

So the total number of different ways is:

OpenStudy (anonymous):

(the product of all those options)

OpenStudy (anonymous):

11

OpenStudy (anonymous):

No... 3*2*9*8*7*6*...*2*1 =

OpenStudy (anonymous):

18144 *2*1

OpenStudy (anonymous):

36288

OpenStudy (anonymous):

I think you're being confused when I say ... I mean the pattern continues.. \[\underbrace{3\cdot2}_{\text{First two songs}} \cdot \underbrace{9\cdot 8\cdot7 \cdot 6 \cdot 5 \cdot 4 \cdot 3 \cdot 2 \cdot 1}_{\text{all the rest}} = 22,177,280\]

OpenStudy (anonymous):

oh ok.. so 3*2= 6* then 9*8*7*6*5*4*3*2*1=

OpenStudy (anonymous):

take the 6 multiply times 9 8 7 6 5 4 3 2 1= my answer

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

i have 2177280 for answer

OpenStudy (anonymous):

i did it 3 times on caculator

OpenStudy (anonymous):

That is what I got also.

OpenStudy (anonymous):

so that would be my answer for the final question

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

ok now i understand this.. Thank you :)